OpenStudy (anonymous):

are 15, 36, and 39 a Pythagorean triple? Show how to prove your answer.

OpenStudy (hoblos):

a,b and c are Pythagorean triple is \[c ^{2} = a^{2} + b^{2}\]

OpenStudy (hoblos):

is 39^2 = 36^2 + 15^2 ?

OpenStudy (anonymous):

see if 15^2+36^2=39^2

OpenStudy (anonymous):

yes it does

OpenStudy (hoblos):

so they are Pythagorean triple